# Learning-based Multi-View Stereo: A Survey

### v1

@@ Abstract
Multi-View Stereo, 3D Reconstruction, Deep Learning.
@@
3D reconstruction describes the general task of recovering the 3D structure of a scene. It is widely applied in augmented/virtual reality (AR/VR), autonomous driving, and robotics [1]. The advancement of 3D acquisition techniques has led to the increased affordability and reliability of depth sensors, such as depth cameras and LiDARs. These sensors are extensively utilized for real-time tasks, enabling rough estimations of the surrounding environment, such as simultaneous localization and mapping (SLAM)  [2, 3, 4, 5] or dense reconstruction  [6, 7, 8]. Nevertheless, depth maps captured by such sensors tend to be partial and sparse, resulting in incomplete 3D representations with limited geometric details. In addition, these sensors are active and usually consume a lot of power. In contrast, camera-based solutions, commonly found in edge devices like smartphones and AR/VR headsets, offer a more economically viable alternative for 3D reconstruction.
@@
One fundamental technique in image-based 3D reconstruction is Multi-View Stereo (MVS). Given a set of calibrated images, MVS aims to reconstruct the dense 3D geometry for an observed scene. Based on the scene representations, traditional MVS methods can be divided into three categories: volumetric, point cloud, and depth map. Volumetric methods [9, 10, 11, 12] discretize the continuous 3D space into voxels and label each as inside or outside of the surface. They are limited to scenes of small scale because of the large memory consumption. Point cloud methods [13, 14] operate directly on 3D points and often employ propagation to gradually densify the reconstruction. As the propagation of point clouds is proceeded sequentially, these methods are difficult to be parallelized and thus take much time in processing [15]. In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes. In contrast, methods relying on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images. Subsequently, these depth maps are fused into a dense representation, e.g., point cloud or mesh, as a post-processing step. Such design decouples the reconstruction problem into per-view depth estimation and depth fusion, which explicitly improves the flexibility and scalability. Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [21, 22, 23, 24].
@@
To overcome these challenges, recent works [15, 25] have turned to learning-based methods, leveraging the emergence of Convolutional Neural Networks (CNNs) that have shown remarkable success in various 2D vision tasks. These learning-based MVS techniques have achieved remarkable success and explicitly outperform traditional methods [18, 17] on various benchmarks [21, 22, 23, 26, 27].
@@
In this survey, we categorize existing learning-based MVS methods based on their characteristics as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Voxel-based methods estimate the geometry with volumetric representation and implicit function, e.g., Signed Distance Functions (SDF). They are limited to small-scale scenes due to the high memory consumption. NeRF and 3D Gaussian Splatting-based methods adapt NeRF [28] and 3D Gaussian Splatting [29], which are originally used for novel view synthesis, to extract surface from the implicit field or point cloud. They typically need to optimize the geometry for each new scene, which needs lots of run-time and memory consumption. Large feed-forward methods typically use large transformer models to directly learn the 3D representation from given images. They require massive computation because of the huge network, and are mainly limited to object-level scenes [30]. Depth map-based methods utilize deep learning in depth estimation and then fuse depth maps with traditional fusion algorithm. Comparatively, depth map-based methods [15, 25] are the main family of learning-based methods since they inherit the advantages from those traditional methods [17, 18, 19, 20] by decoupling 3D reconstruction into depth estimation and depth fusion. Therefore, we focus more on the depth map-based methods and discuss them in more details in this survey.
@@
For clarity, we further categorize depth map-based methods into online and offline methods, shown in Fig. 1. Specifically, online MVS methods [25, 31] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26]. They typically use small resolution inputs and a lightweight network structure to ensure quick inference considering speed and simplicity. The primary focus of online MVS is to provide instant reconstruction for applications where time efficiency is critical, e.g., augmented reality and live video processing. In contrast, offline MVS methods [15, 32] prioritize better quality reconstruction at the cost of computation. These methods excel in multi-view depth estimation and point cloud fusion using high-resolution image sets, e.g., DTU [21], Tanks and Temples [22] and ETH3D [23]. By operating on higher-resolution inputs and employing complex network designs, these approaches thoroughly analyze the multi-view information in the images. The emphasis on achieving high accuracy and capturing intricate details often comes at the cost of requiring considerable computational resources and time. As a result, offline MVS is frequently employed in applications demanding precise and photorealistic scene representations, including 3D modeling and archaeological reconstruction.
@@
In summary, our survey covers the most recent literature on learning-based MVS methods until 2023, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies. Moreover, we summarize the performance of the reviewed methods on different benchmarks, and discuss the potential future directions for deep learning-based MVS methods.
@@
I Preliminaries
Depth map-based MVS, including most traditional and learning-based methods, typically consists of several components: camera calibration, view selection, multi-view depth estimation and depth fusion. In this section, we introduce these components to provide readers a clear picture of the MVS problem. Note that camera calibration and view selection are components for other learning-based methods as well. Sec. I-A introduces camera calibration with Structure from Motion (SfM) or SLAM. Sec. I-B discusses how to select neighboring views to accurately reconstruct the geometry. Sec. I-C explains how to build cost volumes in learning-based MVS methods with plane sweep [33]. Sec. I-D introduces the typical depth fusion strategies after depth estimation. Sec. I-E lists common datasets and benchmarks for MVS and Sec. I-F summarizes the common evaluation metrics.
@@ I Preliminaries
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [34], Furu [14], Gipuma [18]…,leaf,text width=15 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=15 em] ] [Learning-based Methods
with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=25.3em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [31], SimpleRecon [35]…,leaf,text width=18.8 em] ] [Offline Methods [Direct 3D CNN,text width=5 em [MVSNet [15], CIDER [36]…,leaf,text width=12 em] ] [RNN,text width=5 em [R-MVSNet [37], $D^{2}$-RMVSNet [38]…,leaf,text width=12 em] ] [Coarse-to-fine,text width=5 em [CasMVSNet [32], UCSNet [39]…,leaf,text width=12 em] ] [Iterative Update,text width=5 em [PatchmatchNet [40], IterMVS [41]…,leaf,text width=12 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [42], RC-MVSNet [43]…,leaf,text width=10 em] ] [Multi-stage,text width=5 em [U-MVSNet [44], KD-MVS [45]…,leaf,text width=10 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [46],leaf,text width=6.8 em] ] ] [Learning-based Methods
without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [47], NeuralRecon [48]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [49], NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [DreamGaussian [51], SuGaR [52]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [53], DUSt3R [54]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=9.2 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=33.6 em] ] ]
@@ I-B View Selection
I-C Multi-view Depth Estimation with Plane Sweep
Figure 3: Illustration of plane sweep algorithm [33]. To estimate the depth map for reference image (cam_x), neighboring source images (cam_1, cam_2) are projected with homography to fronto-parallel planes of the reference view frustum.
@@ I-C Multi-view Depth Estimation with Plane Sweep
In a nutshell, the plane sweep algorithm in MVS entails iteratively sweeping planes through the object space, computing homographies between images, and selecting depth values based on consensus among different views, ultimately facilitating accurate 3D reconstruction. The plane sweep algorithm discretizes the depth space using a series of parallel planes along the depth direction. It operates by sweeping a conceptual plane through the object space and evaluating the spatial distribution of geometric surfaces. It is worth noting that, for learning-based MVS, plane sweep is conducted purely on GPU, and instead of warping hand-crafted features or edges, the dense pixels are efficiently warped.
We take a toy example of three cameras viewing the same object as illustrated in Fig. 3. Plane sweep is then performed on the frustum of cam_x (termed as the reference camera) by creating a series of fronto-parallel hypothesized planes within a given range (typically $[d_{\textrm{min}},d_{\textrm{max}}]$), with each plane corresponding to a depth value w.r.t. cam_x. Let’s then examine two 3D points, $M$ and $M^{\prime}$, as examples of occupied and unoccupied positions on hypothesized planes, respectively. For $M$, all of the three cameras capture the identical point lying on the geometry surface with photometric consistency. In contrast, the photometric consistency of the observations of $M^{\prime}$ are poor, indicating that $M^{\prime}$ is an invalid hypothesis. To evaluate the similarity of the observations towards $M$ and $M^{\prime}$, the images of cam_1 and cam_2 (termed as the source cameras) are warped to each plane by homography. These warped images are then compared against the scaled images of cam_x. Depth hypotheses with high similarity measures are considered reliable.
@@
The warped source feature maps are then obtained via differentiable interpolation.
Regarding selecting $D$ depth hypotheses from the triangulated depth range $[d_{\textrm{min}},d_{\textrm{max}}]$, there are two main schemes, namely the forward depth sampling and the inverse depth sampling. The naive forward sampling divides the depth range into $D-1$ depth intervals with identical lengths. Given a depth index $k$, we have
$d_{k}=d_{\textrm{min}}+\frac{k}{D-1}(d_{\textrm{max}}-d_{\textrm{min}}),k=0,\ldots,D-1,$ (5)
where the depth hypotheses distribute uniformly between the two ends. While for the inverse sampling scheme [36], we sample uniformly in the multiplicative inverse of $d$, such that
$\frac{1}{d_{k}}=\frac{1}{d_{\textrm{max}}}+\frac{k}{D-1}(\frac{1}{d_{\textrm{min}}}-\frac{1}{d_{\textrm{max}}}),k=0,\ldots,D-1.$ (6)
In this way, the distribution of sampled depth values becomes sparser with $d$ approaching to $d_{\textrm{max}}$. When reconstructing unbounded outdoor scenes, where the depth range is rather large, the inverse sampling will be a reasonable choice since it samples more densely at the foreground. In addition, as revealed in [63], the inverse sampling leads to a uniform sampling on the projected epipolar lines of source images.
@@
So far, for each depth hypothesis made for cam_x, we have the warped source feature maps as well as the scaled reference feature map. The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec. II-B. By performing plane sweep, we obtain a cost volume with a regular spatial shape, making it easy for CNNs to process in parallel.
I-D Depth Fusion
@@
For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh. Online MVS methods [31, 35] usually adopt TSDF (Truncated Signed Distance Function) fusion [64, 65] to fuse the depth maps into a TSDF volume and then use Marching Cube [66] to extract the mesh. However, there usually exist outliers in the depth maps, which may reduce the reconstruction accuracy. To overcome this problem and improve the accuracy, offline MVS methods [15, 32, 40] typically filter the depth maps before fusing into a point cloud, which is motivated by Galliani et al. [18]. There are two main filtering steps: photometric consistency filtering and geometric consistency filtering [15]. For photometric consistency filtering, a per-pixel confidence, discussed in Sec. II-G, is estimated to measure the confidence of depth estimation, i.e., probability that the ground truth depth is within a small range near the estimation. Then a threshold can be set to filter depth values with low confidence. For geometric consistency filtering, the consistency of depth estimations are measured among multiple views. For a pixel $\mathbf{p}$ in the reference view 0, we project it to pixel $\mathbf{p}^{\prime}$ in its $i$-th neighboring view through its depth prediction $d_{0}(\mathbf{p})$. After looking up the depth for $\mathbf{p}^{\prime}$, $d_{i}(\mathbf{p}^{\prime})$, we re-project $\mathbf{p}^{\prime}$ back to the reference view at pixel $\mathbf{p}^{\prime\prime}$ and look up its depth $D_{0}(\mathbf{p}^{\prime\prime})$. We consider pixel $\mathbf{p}$ and its depth as consistent to the $i$-th neighboring view, if the distances, in image space and depth, between the original estimate and its re-projection satisfy:
$\xi_{d}=\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2}\leq\tau_{1},$ (7)
$\xi_{p}=\frac{\|d_{0}(\mathbf{p}^{\prime\prime})-d_{0}(\mathbf{p})\|}{d_{0}(\mathbf{p})}\leq\tau_{2},$ (8)
where $\tau_{1}$ and $\tau_{2}$ are thresholds. The pixels are considered to be reliable estimations if they are consistent in at least $N$ neighboring views. Recently, instead of using predefined thresholds for $\tau_{1}$ and $\tau_{2}$, dynamic consistency checking [38] is proposed to dynamically aggregate geometric matching error among all views and improve the robustness of reconstruction. Specifically, the dynamic multi-view geometric consistency $C_{\textrm{geo}}(\mathbf{p})$ is computed as:
$C_{\textrm{geo}}(\mathbf{p})=\sum_{i}\exp{(-(\xi_{p}+\lambda\xi_{d}))},$ (9)
where $\lambda$ is a weight to balance the reprojection error in two metrics. Then the outliers with $C_{\textrm{geo}}(\mathbf{p})$ smaller than a threshold are filtered.
I-E Datasets and Benchmarks

### v2

@@ Abstract
3D Reconstruction, Multi-View Stereo, Deep Learning.
1 Introduction
@@
3D reconstruction describes the general task of recovering the 3D structure of a scene. It is widely employed in augmented/virtual reality (AR/VR), autonomous driving, and robotics [1]. The advancement of 3D acquisition techniques has led to the increased affordability and reliability of depth sensors, such as depth cameras and LiDARs. These sensors are extensively utilized for real-time tasks, enabling rough estimations of the surrounding environment, such as simultaneous localization and mapping (SLAM)  [2, 3, 4, 5] or dense reconstruction  [6, 7, 8]. Nevertheless, depth maps captured by such sensors tend to be partial and sparse, resulting in incomplete 3D representations with limited geometric details. In addition, these sensors are active and usually consume a lot of power. In contrast, camera-based solutions, commonly found in edge devices like smartphones and AR/VR headsets, offer a more economically viable alternative for 3D reconstruction.
@@
One fundamental technique in image-based 3D reconstruction is Multi-View Stereo (MVS). Given a set of calibrated images, MVS aims to reconstruct dense 3D geometry for an observed scene. Based on the scene representations, traditional MVS methods can be divided into three categories: volumetric, point cloud, and depth map. Volumetric methods [9, 10, 11, 12] discretize the 3D space into voxels and label each as inside or outside of the surface. They are limited to scenes of small scale due to the large memory consumption. Point cloud methods [13, 14] operate directly on 3D points and often employ propagation to gradually densify the reconstruction. As point cloud propagation occurs sequentially, these methods are difficult to parallelize, leading to longer processing times. [15]. In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes. In contrast, methods relying on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images. Subsequently, these depth maps are fused into a dense representation, e.g., point cloud or mesh, as a post-processing step. Such design decouples the reconstruction problem into per-view depth estimation and depth fusion, which explicitly improves the flexibility and scalability. Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [21, 22, 23, 24].
@@
To overcome these challenges, recent works [15, 25] have shifted towards learning-based approaches, have adopted learning-based approaches, using convolutional neural networks (CNNs) that have excelled in 2D vision tasks. These methods have significantly outperformed traditional methods on various benchmarks [21, 22, 23, 26].
@@
In this survey, we categorize existing learning-based MVS methods based on their characteristics as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Voxel-based methods estimate the geometry with volumetric representation and implicit function, e.g., Signed Distance Functions (SDF). They are limited to small-scale scenes due to the high memory consumption. NeRF and 3D Gaussian Splatting-based methods adapt NeRF [27] and 3D Gaussian Splatting [28], which are originally used for novel view synthesis, to extract surface from the implicit field or point cloud. They typically need to optimize the geometry for each new scene, which needs lots of run-time and memory consumption. Large feed-forward methods typically use large transformer models to directly learn the 3D representation from given images. They require massive computation because of the huge network, and are mainly limited to object-level scenes [29]. Depth map-based methods utilize deep learning in depth estimation and then fuse depth maps with traditional fusion algorithm. Comparatively, depth map-based methods [15, 25] are the main family of learning-based methods since they inherit the advantages from those traditional methods [17, 18, 19, 20] by decoupling 3D reconstruction into depth estimation and depth fusion. Therefore, we focus more on the depth map-based methods and discuss them in details.
@@
For clarity, we further categorize depth map-based methods into online and offline methods, shown in Fig. 1. Specifically, online MVS methods [25, 30] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26].These methods typically rely on low-resolution inputs and lightweight network structures to ensure fast inference, prioritizing speed and simplicity. The main goal of online MVS is to deliver real-time reconstruction for time-sensitive applications such as augmented reality and live video processing. In contrast, offline MVS methods [15, 31] prioritize better quality reconstruction at the cost of computation. These methods excel in multi-view depth estimation and point cloud fusion using high-resolution image sets, e.g., DTU [21], Tanks and Temples [22] and ETH3D [23]. By operating on higher-resolution inputs and employing complex network, they thoroughly analyze the multi-view information in the images. The emphasis on achieving high accuracy and capturing intricate details often comes at the cost of requiring considerable computational resources and time. As a result, offline MVS is frequently employed in applications demanding precise and photorealistic scene representations, including 3D modeling and archaeological reconstruction.
@@
In summary, our survey covers the most recent literature on learning-based MVS methods, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies. Moreover, we summarize the performance of the reviewed methods on different benchmarks, and discuss the potential future directions for deep learning-based MVS methods.
@@
2 Preliminaries
Depth map-based MVS, including most traditional and learning-based methods, typically consists of several components: camera calibration, view selection, multi-view depth estimation and depth fusion. In this section, we introduce these components to provide readers a clear picture of the MVS. Note that camera calibration and view selection are components for other learning-based methods as well. Sec. 2.1 introduces camera calibration with Structure from Motion (SfM) or SLAM. Sec. 2.2 discusses how to select neighboring views to reconstruct the geometry. Sec. 2.3 explains how to build cost volumes in learning-based MVS methods with plane sweep [32]. Sec. 2.4 introduces the typical depth fusion strategies after depth estimation. Sec. 2.5 lists common datasets and benchmarks for MVS and Sec. 2.6 summarizes the common evaluation metrics.
@@ 2 Preliminaries
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [33], P-MVS [14], Gipuma [18]…,leaf,text width=16.3 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=16.3 em] ] [Learning-based Methods
with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=27.4 em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [30], SimpleRecon [34]…,leaf,text width=20.2 em] ] [Offline Methods [Direct 3D CNN,text width=5.4 em [MVSNet [15], CIDER [35]…,leaf,text width=13 em] ] [RNN,text width=5.4 em [R-MVSNet [36], $D^{2}$-RMVSNet [37]…,leaf,text width=13 em] ] [Coarse-to-fine,text width=5.4 em [CasMVSNet [31], UCSNet [38]…,leaf,text width=13 em] ] [Iterative Update,text width=5.4 em [PatchmatchNet [39], IterMVS [40]…,leaf,text width=13 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [41], RC-MVSNet [42]…,leaf,text width=10.5 em] ] [Multi-stage,text width=5 em [U-MVSNet [43], KD-MVS [44]…,leaf,text width=10.5 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [45],leaf,text width=6.2 em] ] ] [Learning-based Methods
without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [46], NeuralRecon [47]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [48], NeuS [49], Geo-NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [ SuGaR [51], 2DGS [52], PGSR [53]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [54], DUSt3R [55], MASt3R [56]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=10 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=37.5 em] ] ]
@@ 2.2 View Selection
2.3 Multi-view Depth Estimation with Plane Sweep
@@
In a nutshell, the plane sweep algorithm entails iteratively sweeping planes through the object space, computing homographies between images, and selecting depth values based on consensus among different views, ultimately facilitating accurate 3D reconstruction. The plane sweep algorithm discretizes the depth space using a series of parallel planes along the depth direction. It operates by sweeping a conceptual plane through the object space and evaluating the spatial distribution of geometric surfaces.
@@
The warped source feature maps are then obtained via differentiable interpolation. To better elaborate the process of plane sweep, we include a toy example in the supplementary. The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec. 3.2. By performing plane sweep, we obtain a cost volume with a regular spatial shape, making it easy for CNNs to process in parallel.
2.4 Depth Fusion
@@
For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh. Online MVS methods [30, 34] usually adopt TSDF (Truncated Signed Distance Function) fusion [65, 66] to fuse the depth maps into a TSDF volume and then use Marching Cube [67] to extract the mesh. However, there usually exist outliers in the depth maps, which may reduce the reconstruction accuracy. To overcome this problem and improve the accuracy, offline MVS methods [15, 31, 39] typically filter the depth maps before fusing into a point cloud, which is motivated by Galliani et al. [18]. There are two main filtering steps: photometric consistency filtering and geometric consistency filtering [15]. For more details, please refer to the supplementary.
2.5 Datasets and Benchmarks